OpenStudy (waheguru):

How do I find the x-intercepts of the equation y=-0.25(x+4)^2+1 without graphing

OpenStudy (asnaseer):

what do you think will be the value of y when the curve intersects the x-axes?

OpenStudy (waheguru):

0

OpenStudy (waheguru):

But im getting the wrong answer

OpenStudy (asnaseer):

correct, so, to find the x-intercepts, just set y=0 and solve for x

OpenStudy (asnaseer):

please show your steps so I can spot where you may have made a mistake

OpenStudy (waheguru):

y=-0.25(x+4)^2+1 0=-0.25(x+4)^2+1 -1 = -0.25(x+4)^2 -1/-0.25 = x^2+16

OpenStudy (asnaseer):

(x+4)^2 is not equal to x^2+16

OpenStudy (asnaseer):

also, you do not need to expand this

OpenStudy (waheguru):

So how can I solve from here?

OpenStudy (anonymous):

x + 4 = +-2

OpenStudy (asnaseer):

you got to:\[\frac{-1}{-0.25}=(x+4)^2\]

OpenStudy (asnaseer):

now look at the left-hand-side

OpenStudy (waheguru):

4 = (x+4)^2

OpenStudy (asnaseer):

you have a minus at the top and bottom of the fraction

OpenStudy (anonymous):

\[(x + 4)^2 = 4 \implies x+4 = \pm 2\]

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

then just take square roots of both sides

OpenStudy (waheguru):

Oh, I have to take the square root away first, got you. Thanks

OpenStudy (asnaseer):

that is right - that is why you didn't need to expand (x+4)^2

OpenStudy (waheguru):

can you explain the + or - thing waterineyes did?

OpenStudy (asnaseer):

the square root of 4 is -2 or +2

OpenStudy (asnaseer):

because (-2)^2 = 4 and (2)^2 = 4

OpenStudy (anonymous):

When you take the square root of a number then you put + and - sign in front of it: Because: \[(-x)^2 = x^2\] \[(x)^2= x^2\] From both -x and x you are getting \(x^2\)..

OpenStudy (waheguru):

Thanks alot I never thought I would get the answer

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

So, one value of x you will get by taking +2 and other value you will get by taking -2.. Go ahead,,

OpenStudy (waheguru):

-2-4 = -6 2-4 = -2 This marks the two x-intercepts :)

OpenStudy (waheguru):

sometimes the parabola only has one x-intercept or none how do I know that

OpenStudy (anonymous):

My concepts are not deeper, @asnaseer will help you here..

OpenStudy (asnaseer):

if the equation has two solutions (as you have here), then it has 2 x-intercepts if it has one solution, then there is only one solution if it has no solution then it does not cross the x-axes

OpenStudy (asnaseer):

how much have you studied quadratic equations? have you heard about the determinant?

OpenStudy (waheguru):

no, not yet

OpenStudy (waheguru):

I guess I will find out later on then

OpenStudy (asnaseer):

I think you will. The general quadratic has the form:\[ax^2+bx+c=0\]it's determinant is given by:\[b^2-4ac\]where a, b and c are constants

OpenStudy (asnaseer):

if the determinant is positive, then the quadratic has 2 solutions if the determinant is zero, then the quadratic has 1 solution and if the determinant is negative, then the quadratic has no solutions

OpenStudy (asnaseer):

you can visualise this as follows:

OpenStudy (asnaseer):